OpenStudy (anonymous):

Trig / Pre Cal/ identities Guide me please. \( \frac{\frac{\sin x}{\cos x}}{\frac{1}{cos x} + 1 \)

OpenStudy (anonymous):

It is suppose to be. \( \frac{\frac{\sin x}{\cos x}}{\frac{1}{cos x}+1 }= \frac{\sec x+1}{\tan x} \)

OpenStudy (anonymous):

\( \large \rm \frac{\frac{\sin x}{\cos x}}{\frac{1}{cos x}+1 }= \frac{\sec x+1}{\tan x} \)

OpenStudy (anonymous):

These things are killing me lol

OpenStudy (anonymous):

not an identity

OpenStudy (xapproachesinfinity):

okay so we have \[\Huge \frac{\frac{\sin x}{\cos x}}{\frac{1}{\cos x}+1}=\frac{\sec x+1}{\tan x }\]

OpenStudy (anonymous):

I got that they're reciprocals

OpenStudy (xapproachesinfinity):

if so we can do this \[\Huge \frac{\tan x}{\sec x+1}=\frac{\sec x+1}{\tan x}\] clearly this is not truee

OpenStudy (xapproachesinfinity):

try some vlues if you have doubt

OpenStudy (anonymous):

Yes but I am trying to get from the left side to the right. \(\Huge \frac{\frac{\sin x}{\cos x}}{\frac{1}{\cos x}+1}=\frac{\sec x+1}{\tan x }\) Do I times by cos?? \(\Huge \frac{\frac{\sin x}{\cos x}}{\frac{1}{\cos x}+1} \times cos x\)

OpenStudy (anonymous):

The identity is false is what he's saying. You won't be able to get from the left to the right.

OpenStudy (xapproachesinfinity):

now it does not work! left does not equal right

OpenStudy (xapproachesinfinity):

plug in some values to be sure if you have doubt

OpenStudy (xapproachesinfinity):

no* for (now)

OpenStudy (anonymous):

The original problem is \(\Huge \frac{tan x}{\sec - 1 }=\frac{\sec x+1}{\tan x }\)

OpenStudy (xapproachesinfinity):

0 for example is not good left give you 0 right gives you undefined 2/0 if it is an identity it should not be that way

OpenStudy (anonymous):

I did the left side assuming that is what we have to do.

OpenStudy (xapproachesinfinity):

is the left 1/cosx -1 or 1/cosx +1 you just changed in your last reply?

OpenStudy (xapproachesinfinity):

in that case it is an identity

OpenStudy (xapproachesinfinity):

see that you give us the wrong thing from the start

OpenStudy (xapproachesinfinity):

if it is a minus on the bottom it is an identity since tan^2x=sec^2x-1

OpenStudy (anonymous):

This is the original problem. \(\Huge \frac{tan x}{\sec - 1 }=\frac{\sec x+1}{\tan x } \) How do we know to work the numerator or the denominator or both?

OpenStudy (anonymous):

For instance, should I \(\Huge \frac{tan x}{\frac{1}{cos x}-1 }=\frac{\sec x+1}{\tan x } \) OR \(\Huge \frac{\frac{sin}{cos}}{{\sec x}-1 }=\frac{\sec x+1}{\tan x } \) OR \(\Huge \frac{\frac{\sin x}{\cos x}}{\frac{1}{\cos x}+1} \)

OpenStudy (xapproachesinfinity):

usually when you prove something you should get from left to right of vice versa so taking left \[\frac{\tan x}{\sec x=1}=\frac{\tan x(\sec x+1)}{(\sec x-1)(\sec x+1)}\] \[=\frac{\tan x(\sec x+1)}{\tan^2x }=\frac{\sec x+1}{\tan x} ~~~Q.E.D\]

OpenStudy (xapproachesinfinity):

you didn't need to do all that just leave tan and sec that way

OpenStudy (xapproachesinfinity):

you are making it hard

OpenStudy (anonymous):

There are no rules though

OpenStudy (xapproachesinfinity):

what rules?

OpenStudy (xapproachesinfinity):

did you get what i did?

OpenStudy (anonymous):

No

OpenStudy (xapproachesinfinity):

if we do it your way we have \[\frac{\sin x}{1-\cos x}=\frac{\sin x(1+\cos x)}{1-\cos^2x}=\frac{1+\cos x}{\sin x}\] multiplyinh by 1/cos top and bottom we get \[=\frac{\sec x+1}{\tan x}\]

OpenStudy (xapproachesinfinity):

perhaps i'm going fast?

OpenStudy (xapproachesinfinity):

i skip some stuff given that you know how to deal with algebra

OpenStudy (anonymous):

I don't get it. How do I know to work the numerator or the denominator or both. How do I know what method to use when there are no rules to guy you.

OpenStudy (anonymous):

For instance, how did you do it your way?

OpenStudy (xapproachesinfinity):

like \[\huge \frac{\frac{a}{b}}{\frac{1}{b}-1}=\frac{a}{b(\frac{1}{b}-1)}=\frac{a}{1-b}\]

OpenStudy (xapproachesinfinity):

this is what i did with your way changing tan to sin/cos and sec to 1/cos

OpenStudy (anonymous):

Yes that is my way.

OpenStudy (anonymous):

But there are other ways and some ways don't work

OpenStudy (xapproachesinfinity):

yes once you get sin x/1-cosx you multiply by 1+cosx top and bottom

OpenStudy (anonymous):

Correct

OpenStudy (xapproachesinfinity):

what other Ways do you mean if you are using identities they should work

OpenStudy (xapproachesinfinity):

my way is no different than yours i still had to multiply top and bottom be 1+sec x to get sec^2x-1 on the bottom which is the same as tan^2x

OpenStudy (xapproachesinfinity):

this is just a matter of knowing your identities and how to use them

OpenStudy (anonymous):

For instance \( \Huge \frac{tan x}{\frac{1}{cos x}-1 }=\frac{\sec x+1}{\tan x } \) \( \Huge \frac{(cosx)(tan x)}{cos x(\frac{1}{cos x}-1) }=\frac{\sec x+1}{\tan x } \)

OpenStudy (xapproachesinfinity):

it is still working since top becomes sinx

OpenStudy (xapproachesinfinity):

gotta go! i think you got this now?

OpenStudy (anonymous):

Ok so we have \(\Huge \frac{(cosx)(tan x)}{cos x(\frac{1}{cos x}-1) }=\frac{\sec x+1}{\tan x } \) \(\Huge \frac{sin x}{1-cos x}=\frac{\sec x+1}{\tan x } \)

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (xapproachesinfinity):

i explained above how you obtain the right side